Calculate the volume of a sphere with a diameter of 10 inches. Round answer to the nearest cubic inch.
Answer by nerdybill(5404)

Calculate the volume of a sphere with a diameter of 10 inches. Round answer to the nearest cubic inch.
.
Volume of a sphere = (4/3)(pi)r^3
where
pi is 3.14
r is radius
.
radius = (1/2)diameter = (1/2)10 = 5 inches
.
Volume of a sphere = (4/3)(pi)r^3
Volume of a sphere = (4/3)(3.14)5^3
Volume of a sphere = (4/3)(3.14)125
Volume of a sphere = (166.667)(3.14)
Volume of a sphere = 523.3333
Volume of a sphere = 523 cubic inches